Operations Report – April 07th

Crew 191 Operations Report 07/04/2018

SOL: 14

Name of person filing report: Wataru Okamoto

Non-nominal systems: Water Tank

Notes on non-nominal systems: We cleaned up all tank. Loft water filter is already ordered by Dr. Shannon. Now water line is not filter line,another line bypassed.

Generator (hours run): 12hours

Solar SOC 83% (Before generator is run at night)

Diesel 80%

Propane 70psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? gallons.

Water (trailer) 0%

Water (static) 60%

Trailer to Static Pump used ? Yes

Water (loft) Static to Loft Pump used Yes

Water Meter: 133208.1
Toilet tank emptied: No

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used:
Hours the ATVs were Used today:

Notes on ATVs:

Deimos rover used: Yes
Hours: 118.6
Beginning charge: %
Ending charge: %
Currently charging: plugged

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 30.2h
Beginning charge: 100%
Ending charge: 100%
Currently charging:

Opportunity rover used: –
Hours: 28.8h
Beginning charge: 100%
Ending charge: 100%
Currently charging:

Curiosity rover used: Yes

Hours: 25.1h
Beginning charge: 100%
Ending charge: 100%
Currently charging: plugged

HabCar used and why, where? We went to Hanksville 3 times. Because water supply.

General notes and comments: We explain to CREW192 about engineering.

Summary of internet: No problem

Summary of suits and radios: 5 charger of the backpack out of order.

Summary of Hab operations: We supply the black tank 20%.

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: CREW192 member will sleep inside the dome tonight.

Summary of RAMM operations: N/A

Summary of health and safety issues: No Problem

Questions, concerns and requests to Mission Support: Thank you for your support. We completed the mission.

Greenhab Report – April 07th

GreenHab Report
Kai Takeda
7 APR 2018

Environmental control: (Choose which of the following is appropriate and explain further if needed)

Ambient with window/door open
Shade cloth on
Fan on(11:00~16:15)

Average temperatures: 25.4℃
Low 14.6℃
High 41.7℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 4gal

Time(s) of watering for crops: twice

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

I planted our seedling on pots in GreenHab.

Support/supplies needed: None

Crew Photos – April 07th

SOL Essay (Photo Description)
07 Apr 2018

Sol 14
Miho Tsukishiro (XO)

In this morning, we did an additional EVA for Japanese TV broadcasting.
Meanwhile, the rest of crew worked hard with clean up and packing our staff (aibo too!) for check out the HAB. And also, we are pleased to announce that: We finally cleaned the loft tank. That is, the water tank problem in HAB has solved completely!
Before we took the things in HAB over to the following crew 192, we went on a small trip on the Hab Ridge. How amazing the landscape is.
Now we felt great satisfaction with coming here MDRS. Thank you Mission Support Team again!

Sol Summary – April 07th

Crew 191 Sol Summary Report
07 Apr 2018

Sol 14

Summary Title:
Cleaning and taking the HAB over to following crew
Miho Tsukishiro (XO)

Mission Status:

Sol Activity Summary:
7:20 Breakfast
7:45 Briefing
9:00 EVA #17 The additional EVA (just for TV broadcasting)
13:00 Welcoming crew192
13:30 Lunch
14:00 Cleaning & Transportation of Water from Hollow Mountain
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
leaving MDRS

Anomalies in work:

Cloudy. Mild winds.

Crew Physical Status:


Reports to be file:
– SOL Summary Report
– Engineer Report
– GreenHab Report
– Photo Report with Description

Support Requested:

Copyright © The Mars Society. All rights reserved. | Main Site